The sum of infinite geometric series with first term 3 and second term 0.75 is?
Answer by htmentor(1343)

The sum of the series is given by:
S = a/(1-r) where a = the 1st term in the series and r = the common ratio
r = 0.75/3 = 0.25
So S= 3/0.75 = 4